A roof rises 5.5 ft over a horizontal distance of 18.58 ft. What is the slope of the roof to the nearest hundredth?
Answer by edjones(8007)

Let one point be (0,0)
let the other be (18.58,5.5)
slope=change in y/change in x=rise/run=5.5-0/18.58-0=5.5/18.58=.30
Ed